OpenStudy (anonymous):

Solve ln (2x + 3) = 7.

OpenStudy (anonymous):

(2x + 3) = exp(7) 2x = exp(7) -3 solve for x

OpenStudy (anonymous):

(2x+3)=7 2x+3=7 2x+3-3=7-3 2x=4 x=4/2 x=2 Answer= 2 :)

OpenStudy (anonymous):

What happened to the logarithm and the exponential function?